A put option and a call option on a stock have the same expiration date and the same exercise (or strike price). Both options expire in 6 months. Assume that put-call parity holds and interest rate is positive. If both call and put options have the same price, which of the following is true?

A) Put option is in-the-money.

B) Call option is in-the-money.

C) Both call and put options are in-the-money.

D) Both call and put options are out-of-the-money.
Answer #1

Option A

According to Put Call parity
S+P=C+K/(1+r)^t

If P=C, S=K/(1+r)^t

i.e, Stock price is less than strike price hence put is in the money and call is out of the money